OpenStudy (tomektomeknyc):

cos (2 arc sin 1) = a) 2 pi b) 1 c) 0 d) - 1

OpenStudy (anonymous):

\[ \cos( 2 \arcsin (1))=\cos\left (2 \frac \pi 2\right )=\cos(\pi) \]

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

What is \[ \cos(\pi) \]

OpenStudy (tomektomeknyc):

-1

OpenStudy (anonymous):

Great

OpenStudy (tomektomeknyc):

thank you very much

OpenStudy (anonymous):

yw

OpenStudy (tomektomeknyc):

how did you know it is cos 2 pi/2 ?

OpenStudy (tomektomeknyc):

are there any tutorials or formulas that show it?

OpenStudy (anonymous):

Which angle has sin=1?

OpenStudy (anonymous):

Try using my site on http://www.math.missouri.edu/escgi/mucgi-bin/munew.cgi?variable=trig

OpenStudy (tomektomeknyc):

thank you